OpenStudy (anonymous):

3/x-2+21/x^2-4=14/x+2 helpppppppppppp meeeeeeeeeeeeee plzzzzzzzzzz

OpenStudy (anonymous):

help me plz

OpenStudy (anonymous):

help me plz

OpenStudy (math&ing001):

\[\frac{ 3 }{ x-2 }+\frac{ 21 }{ x ^{2}-2^{2} }=\frac{ 3(x+2)+21 }{ (x-2)(x+2) }=\frac{ 3x+27 }{ (x-2)(x+2) }=\frac{ 14 }{ x+2 }\]

OpenStudy (anonymous):

yes\

OpenStudy (anonymous):

is thats it or

OpenStudy (math&ing001):

\[3x+27=14(x-2) \] solve for x.

OpenStudy (anonymous):

i got 5

OpenStudy (anonymous):

than how do i get solution set

OpenStudy (anonymous):

u their

OpenStudy (math&ing001):

Your answer is correct. Gratz! ^^

OpenStudy (anonymous):

but how do i find the solution set

OpenStudy (math&ing001):

You only got one solution here which is 5. \[solution set= S =\left\{ 5 \right\}\]

OpenStudy (anonymous):

k thanks